A production facility contains two machines that are used to rework items that are initially defective. Let 𝑋 be the number of hours that the first machine is in use and let π‘Œ be the number of hours that the second machine is in use, on a randomly chosen day. Assume that 𝑋 and π‘Œ have a joint probability density function given by 𝑓(π‘₯) = { 3 2 (π‘₯ 2 + 𝑦 2 ) 0 < π‘₯ < 1 π‘Žπ‘›π‘‘ 0 < 𝑦 < 1 0 π‘œπ‘‘β„Žπ‘’π‘Ÿπ‘€π‘–π‘ π‘’. a. What is the probability that both machines are in operation for less than half an hour?Β 


1
Expert's answer
2022-01-12T18:11:31-0500

"P(0<x<0.5,0<y<0.5)=\int^{0.5}_0\int^{0.5}_0 f(x)dxdy="


"=\int^{0.5}_0\int^{0.5}_0 \frac{3}{2}(x^2+y^2)dxdy=\frac{3}{2}\int^{0.5}_0(x^3/3+xy^2)^{0.5}_0dy="


"=\frac{3}{2}\int^{0.5}_0(1/24+y^2/2)dy=\frac{3}{2}(y/24+y^3/6)|^{0.5}_0="


"=\frac{3}{2}(1/48+1/48)=1/16"